# Benchmarking Large Language Models for Cryptanalysis and Mismatched-Generalization

### v1

@@ A.4 Dataset Sample and Statistics
A.5 Partial Comprehension
The Table 10 shows some examples of the results of the decryption with good comprehension but fragile decryption. In the first example, the decryption is largely accurate, with the only error being the substitution of "patients" with "patience." This suggests strong overall comprehension, but minor challenges in precise lexical replication. In the sixth example, although the model successfully reconstructs the sentence structure, it fails to decrypt a single critical word. Additionally, the fifth example exhibits a substitution error in which a name is altered, indicating potential weaknesses in handling proper nouns and specific identifiers.
Model Plain Text Cipher Decrypted Text NL
A.6 Decryption Difficulty Analysis
@@
Rail Fence $O(n)$ $n-1$ Medium
@@
Playfair $O(n)$ $26!$ Medium
RSA $O(n^{3})$ Large num. Hard
AES $O(n)$ $2^{128}$ Hard
Table 11: Encryption Algorithms Analysis with n as text length Complexity
Refering to Table 11, the key space is the set of all valid, possible, distinct keys of a given cryptosystem. Easy algorithms, such as the Caesar Cipher (key space: 26 for English alphabet), Atbash (key space: 1, fixed mapping by alphabet reversal), and Morse Code (no key, we use standard morse encoding) are classified as trivial to decrypt due to their limited key spaces and straightforward implementation. These algorithms have a linear time complexity of $O(n)$ for both encryption and decryption, making them highly susceptible to brute-force attacks and frequency analysis. The Bacon cipher, despite its binary encoding nature, also falls into this category with its fixed substitution pattern.
@@
Similarly, Playfair cipher (Medium) uses a 5x5 key grid setup resulting in a substantial key space of $26!$ possible arrangements. Its operational complexity is $O(n)$ for both encryption and decryption as each character pair requires only constant-time matrix lookups. Playfair is classified as medium due to its resistance to simple frequency analysis and the computational effort required for key search (i.e. 26! arrangements).
RSA (Hard) is a public-key encryption algorithm that relies on the mathematical difficulty of factoring large numbers. Its complexity is $O(n^{3})$ due to the modular exponentiation involved in encryption and decryption. The security of RSA comes from its large key space and the computational infeasibility of breaking it without the private key.
While AES (Hard) has an $O(n)$ time complexity for encryption/decryption operations, its security derives from an enormous key space ($2^{128}$, $2^{192}$, or $2^{256}$, depending on key size) combined with sophisticated mathematical properties that make cryptanalysis computationally infeasible. In addition, AES’s security also depends on its round-based structure and strong avalanche effect, making it resistant to both classical and modern cryptanalytic attacks.
A.7 Evaluating Metrics
@@
Normalized Levenshtein Yujian and Bo (2007) is used for a more nuanced character-level evaluation which also accounts for the order of characters. To enhance interpretability, we employ a formalized version of this metric, the Levenshtein Decision, defined as:
Normalized $\text{Normalized Levenshtein}=\frac{L(s_{1},s_{2})}{\max(\text{len}(s_{1}),\text{len}(s_{2}))}$
@@
A.8 Cipher Classification
We prompt the LLMs to hypothesize which encryption method was utilized, based solely on the provided ciphertext. This is crucial because if LLMs can identify encryption methods without training, it might enable more sophisticated evasion techniques in malicious prompts, posing significant security risks in sensitive applications. We do not use a separate prompt but in combination with our decryption prompts A.3. We note that this is not about classification in the traditional sense, but rather about assessing the models’ comprehension and interpretative skills when faced with encrypted data. The score improvements after few-shot reflect the models’ ability to identify ciphers from a single-shot example. This improvement is noteworthy as it can be used for jailbreaking attacks by providing obfuscation details or few-shot examples as a context (ICL). Table 6 presents a comparative analysis of six large language models’ performance in zero-shot classification of encryption algorithms.
@@ A.8 Cipher Classification
A.9 Tokenization Inflation Issues in Encrypted Texts
Our token analysis reveals a dramatic token distribution shift post-encryption (13.66× for RSA, 7.93× for Bacon, 6.90× for Morse), exposing two distinct failure modes. While RSA’s security holds cryptographically, Bacon and Morse (Easy) - diverge sharply in decipherment success presumably due to pretraining exposure differences. Similar to Caesar cipher, Morse code benefits from abundant pretraining data (".-" patterns appear frequently in pre-training texts), enabling models to learn dot-dash mappings despite 6.9× token inflation.
Cipher Avg. Token Length Ratio to Plaintext
Normal Text 95.86 1.00x
Caesar Cipher 237.72 2.48x
Atbash Cipher 233.97 2.44x
Morse Code 661.39 6.90x
Bacon Cipher 760.36 7.93x
Playfair Cipher 218.04 2.27x
Rail Fence Cipher 218.64 2.28x
Vigenère Cipher 230.97 2.41x
RSA Cipher 1309.00 13.66x
AES Cipher 457.08 4.77x
Table 12: Comparison of cipher token lengths relative to plaintext
The Atbash cipher (Easy) showed limited comprehension despite low pre-training data, thanks to generalization. In contrast, the Bacon cipher (Easy) completely failed because its rare AB combinations (AB, ABB, ABA) suffer from catastrophic tokenization—expanding 7.93× in length and losing structure. Unlike Atbash’s learnable patterns, Bacon’s repetitive sequences are misinterpreted as character repeats, leading to a distribution shift that prevents meaningful generalization.
A.10 Other Tables and Figures
Figure 3: Distribution of Length by Text Type
@@ A.10 Other Tables and Figures
Table 15: Performance Across styles of writing with focus on Exact Match (EM) and Normalised Levenshtein. Here, Normal Text represents average score for all other types of text.

### v2

@@ A.3 Dataset and Statistics
A.4 Decryption Difficulty Analysis
@@
Rail Fence $O(n)$ n1n-1 Medium
@@
Playfair $O(n)$ $25!$ Medium
RSA $O(k^{3})$ or GNFS Large num. Hard
AES $2^{n/2}$ $2^{128}$ Hard
Table 12: Encryption Algorithms Analysis with n as text length Complexity.
Referring to Table 12, the key space is the set of all valid, possible, distinct keys of a given cryptosystem. Easy algorithms, such as the Caesar Cipher (key space: 26 for English alphabet), Atbash (key space: 1, fixed mapping by alphabet reversal), and Morse Code (no key, we use standard morse encoding) are classified as trivial to decrypt due to their limited key spaces and straightforward implementation. These algorithms have a linear time complexity of $O(n)$ for both encryption and decryption, making them highly susceptible to brute-force attacks and frequency analysis. The Bacon cipher, despite its binary encoding nature, also falls into this category with its fixed substitution pattern.
@@
Similarly, Playfair cipher (Medium) uses a 5x5 key grid setup resulting in a substantial key space of $25!$ possible arrangements. Its operational complexity is $O(n)$ for both encryption and decryption as each character pair requires only constant-time matrix lookups. Playfair is classified as medium due to its resistance to simple frequency analysis and the computational effort required for key search (i.e. 25! arrangements).
RSA (Hard) is a public-key encryption algorithm that relies on the mathematical difficulty of factoring large numbers. Considering runtime brute-force, its complexity is $O(n^{3})$ due to the modular exponentiation involved in encryption and decryption. However, breaking RSA (recovering the private key) reduces to integer factoring, for which the best classical attacks Briggs (1998)’s the General Number Field Sieve (GNFS) run in sub‑exponential time in the modulus size rather than polynomial time. In practice RSA security is therefore expressed in terms of modulus bit length (e.g., 2048 bits) and the infeasibility of known factoring methods for those sizes.
AES (Hard) is a symmetric block cipher whose encryption/decryption cost is linear in the number of 128‑bit blocks processed (i.e., per‑block operations are constant time), so operational runtime is essentially proportional to message length. Its cryptographic strength comes from the large brute‑force key spaces ($2^{128}$, $2^{192}$, or $2^{256}$ for AES‑128/192/256) together with design properties (round structure, diffusion/confusion, and resistance to known structural attacks) that make practical cryptanalysis infeasible. Note also that generic quantum search, termed Grover’s algorithm Grover (1996) would only square‑root the key‑search cost, reducing a $2^{n}$ brute‑force effort to roughly $2^{n/2}$, which is why larger symmetric keys are recommended in post‑quantum planning.
A.5 Evaluating Metrics
@@
Normalized Levenshtein Similarity Yujian and Bo (2007) is used for a more nuanced character-level evaluation which also accounts for the order of characters. To enhance interpretability, we employ a formalized version of this metric, the Levenshtein Similarity, defined as:
$\text{NL}=1-\frac{L(s_{1},s_{2})}{\max(\text{len}(s_{1}),\text{len}(s_{2}))}$
@@
A.6 Random‑guessing baseline (“Lorem Ipsum”)
To better contextualize non‑zero NL Similarity values for apparently incorrect outputs, we evaluated a random‑guessing baseline (“Lorem Ipsum”) to estimate metric bias. The random baseline yields EM $\simeq$ 0 and BLEU $\approx$ 0 (as expected), but a consistently non‑zero NL (similarity) of $\approx$ 0.18–0.19 across all ciphers, indicating a positive bias for NL on arbitrary text. Consequently, we (i) treat NL values near $\sim$0.18 as the floor for “completely incorrect” guesses and (ii) rely primarily on EM and BLEU for strict correctness and NL (above the random baseline) for graded similarity.
Cipher (complexity) EM BLEU NL
Caesar (Easy) 0.00 0.0004 0.18
Atbash (Easy) 0.00 0.0004 0.19
Morse (Easy) 0.00 0.0004 0.18
Bacon (Easy) 0.00 0.0004 0.19
Rail F. (Medium) 0.00 0.0004 0.19
Playfair (Medium) 0.00 0.0004 0.19
Vigenere (Medium) 0.00 0.0004 0.18
AES (Hard) 0.00 0.0004 0.18
RSA (Hard) 0.00 0.0004 0.18
Overall 0.00 0.0004 0.18
Table 13: Random‑guessing baseline results (EM, BLEU, Normalized Levenshtein).
The inclusion of a random-guessing baseline demonstrates that NL scores exhibit a consistent positive bias ($\sim$0.18–0.19) for random outputs, across all encryption schemes; whereas BLEU and EM scores remain robust (near-zero for random text).
A.7 Cipher Classification
We prompt the LLMs to hypothesize which encryption method was utilized, based solely on the provided ciphertext. This is crucial because if LLMs can identify encryption methods without training, it might enable more sophisticated evasion techniques in malicious prompts, posing significant security risks in sensitive applications. We do not use a separate prompt but in combination with our decryption prompts A.2. We note that this is not about classification in the traditional sense, but rather about assessing the models’ comprehension and interpretative skills when faced with encrypted data. The score improvements after few-shot reflect the models’ ability to identify ciphers from a single-shot example. This improvement is noteworthy as it can be used for jailbreaking attacks by providing obfuscation details or few-shot examples as a context (ICL).
@@ A.7 Cipher Classification
A.8 Text Length Additional Analysis
In Table 14, we evaluate decryption on ciphers having varying token lengths.
Model EM NL
Claude-3.5 Sonnet
30 char 0.42 0.55
100 char 0.41 0.53
$\sim$500 tokens 0.36 0.37
GPT-4o
30 char 0.29 0.46
100 char 0.19 0.35
$\sim$500 tokens 0.08 0.16
Table 14: Decryption performance across different sequence lengths.
We note that cryptanalytic capabilities deteriorate significantly with longer sequences, especially for GPT line of models. For Claude-3.5 Sonnet, EM performance drops from 0.42 to 0.37 (12% decrease) and NL drops from 0.55 to 0.36 (35% decrease) when moving from 30 characters to  500 tokens. GPT-4o shows drastic degradation, with EM dropping from 0.29 to 0.08 (72% decrease) and NL from 0.46 to 0.16 (65% decrease).
A.9 Other Tables and Figures
@@
Table 15: Performance Across styles of writing with focus on Exact Match (EM) and Normalised Levenshtein Similarity. Here, Normal Text represents average score for all other types of text.